$SPAD/src/input ipftest.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool ipftest.output
)set message test on
)set message auto off
)clear all
--S 1 of 8
gf2 := PF 2
--R
--R
--R   (1)  PrimeField(2)
--R                                                                 Type: Domain
--E 1
--S 2 of 8
a : gf2 := primitiveElement()$gf2
--R
--R
--R   (2)  1
--R                                                          Type: PrimeField(2)
--E 2
--S 3 of 8
order a
--R
--R
--R   (3)  1
--R                                                        Type: PositiveInteger
--E 3
--S 4 of 8
primitive? a
--R
--R
--R   (4)  true
--R                                                                Type: Boolean
--E 4
--S 5 of 8
createPrimitivePoly(2)$FFPOLY(gf2)
--R

--R
--R         2
--R   (5)  ?  + ? + 1
--R                              Type: SparseUnivariatePolynomial(PrimeField(2))
--E 5
--S 6 of 8
createPrimitivePoly(4)$FFPOLY(gf2)
--R
--R
--R         4
--R   (6)  ?  + ? + 1
--R                              Type: SparseUnivariatePolynomial(PrimeField(2))
--E 6
--S 7 of 8
createPrimitivePoly(12)$FFPOLY(gf2)
--R
--R
--R         12    6    4
--R   (7)  ?   + ?  + ?  + ? + 1
--R                              Type: SparseUnivariatePolynomial(PrimeField(2))
--E 7
--S 8 of 8
createPrimitivePoly(5)$FFPOLY(PF 3)
--R
--R
--R         5    3
--R   (8)  ?  + ?  + ? + 1
--R                              Type: SparseUnivariatePolynomial(PrimeField(3))
--E 8
)spool
)lisp (bye)

———-

References^{[1] nothing}

5